OpenStudy (jazzyluvsu123):

x/6=7.5 I know I'm supposed to multiply but am I multiplying by 6?

OpenStudy (jazzyluvsu123):

@mathstudent55 @satellite73

OpenStudy (4everaddicted2anime):

yes

OpenStudy (jazzyluvsu123):

so 7.5*6? my answer would be 45

OpenStudy (anonymous):

Step 1: Multiply both sides by 6. x/6=7.5

OpenStudy (anonymous):

(x/6)*(6)=(7.5)*(6)

OpenStudy (anonymous):

Answer: x=45